OpenStudy (ronaldo7):

If f(x)=3x^2+4x+5 Evaluate f(x+h)-f(x)/h

OpenStudy (ronaldo7):

help

OpenStudy (jiteshmeghwal9):

F(x+h)=3(x+h)^2+4(x+h)+5 F(x)=3x^2+4x+5

OpenStudy (jiteshmeghwal9):

Find F(x+h)-f(x) ----------------- h

OpenStudy (ronaldo7):

im confused

OpenStudy (ronaldo7):

im getting a different answer when i solve

OpenStudy (jiteshmeghwal9):

\[\frac{3(x+h)^2-3x^2+4(x+h)-4x+5-5}{h}\]

OpenStudy (ronaldo7):

you mean distribute the negative through f(x)

OpenStudy (jiteshmeghwal9):

\[\frac{3(x^2+2hx+h^2)-3x^2+4(x+h)-4x}{h}\]

OpenStudy (jiteshmeghwal9):

Yes i mean exactly that

OpenStudy (ronaldo7):

the answer by my teacher is 6x+3h+4 im not getting that

OpenStudy (ronaldo7):

hello?

OpenStudy (jiteshmeghwal9):

\[\frac{3x^2-3x^2+6hx+3h^2+4x-4x+4h}{h}\]

OpenStudy (ronaldo7):

oh hold on

OpenStudy (ronaldo7):

got it now thank you

OpenStudy (jiteshmeghwal9):

Yw